High Desert Potteryworks makes a variety of pottery products that It sells to retallers. The company uses a job-order costing system In which departmental predetermined overhead rates are used to apply manufacturing overhead cost to Jobs. The predetermined overhead rate In the Molding Department is based on machine-hours, and the rate In the Palnting Department is based on direct labor hours. At the beginning of the year, the company provided the following estimates: Direct labor-hours Machine-hours Fixed manufacturing overhead cost Variable manufacturing overhead per machine-hour Variable manufacturing overhead per direct labor-hour Department Molding Painting 36,588 58,5ee 89,888 31,888 $178,88 $514,888 $ 2.88 4.88 Job 205 was started on August 1 and completed on August 10. The companys cost records show the following Information concerning the job: Department MoldingPainting Direct labor-hours Machine-hours Direct materials Direct labor cost 125 75 $ 944 $1,288 $ 758 $ 988 83 388 Required 1. Compute the predetermined overhead rates used In the Molding Department and the Palnting Department. 2. Compute the total overhead cost applied to Job 205 3-a. What would be the total manufacturing cost recorded for Job 205? 3-b. If the Job contalned 27 units, what would be the unit product cost?
Answer #1

1. Predetermined overhead rates used in the Molding Department:

Fixed manufacturing overhead per machine hour = Fixed manufacturing overhead cost / machine-hours

= $178,000 / 89,000 = $ 2

Variable manufacturing overhead per machine hour = $ 2.80

Predetermined overhead rate = $2 + $2.80 = $4.80

Predetermined overhead rate used in the Painting Department:

Fixed manufacturing overhead per direct-labor hour = Fixed manufacturing overhead cost / direct-labor hours

= $514,800 / 58,500 = $ 8.80

Variable manufacturing overhead per direct-labor hour = $ 4.80

Predetermined overhead rate = $8.80 + $4.80 = $13.60

2. Total overhead cost applied to Job 205:

Molding Department = 380 machine hours x $4.80 = $ 1,824

Painting Department = 125 direct-labor hours x $13.60 = $ 1,700

Total overhead cost = $ 1,824 + $ 1,700 = $ 3,524

3-a. Total manufacturing cost recorded for Job 205:

Molding Department = Direct materials + Direct labor cost + Overhead cost = $944 + $750 + $1,824 = $3,518

Painting Department = Direct materials + Direct labor cost + Overhead cost = $1,280 + $980 + $1,700 = $3,960

Total manufacturing cost = $3,518 + $3,960 = $ 7,478

3-b. Unit product cost = $ 7,478 / 27 units = $ 276.96